If events A and B are mutually exclusive and the P(A) = 3 over 4 and the P(B) = 1 over 8 , what would be the P(A or B)?

Answer by fractalier(6550)

P(A or B) = P(A) + P(B) - P(A and B) = 3/4 + 1/8 - (3/4)(1/8) = 7/8 - 3/32 =
25/32